OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 36? A) x squared over 36 B) ±6square root of x C) 1 over quantity x squared minus 36 D) ±square root of the quantity x plus 36

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

f(x) = x^2 - 36 y = x^2 - 36 .... replace f(x) with y x = y^2 - 36 .... swap x and y now solve for y to get the inverse